OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (anonymous):

-4x - 9y = -19 -10y = -30 ----- 4x + 3y = 5 2y = -6 ----- 7x - 3y = -11 9x = -6 ----- 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

wanna go back to the twiddla or just work here?

pooja195 (pooja195):

solve all the problems now that u know how to :)

OpenStudy (anonymous):

i'm trying but can't get nothing

OpenStudy (anonymous):

i starting with 4x - y = -11

OpenStudy (anonymous):

ooh wait i am doing something else give me a sec

OpenStudy (anonymous):

it's A, rught @pooja195

OpenStudy (anonymous):

come when you are ready :) @pooja195

pooja195 (pooja195):

A is right :)

OpenStudy (anonymous):

thanks :)